OpenStudy (anonymous):

find the sum of the infinite geometric series, if possible. 8+6+9/2+27/8........ and the book says the answer it 32. Explain, anybody?? I don't understand what pattern or rule this set of numbers is following....

OpenStudy (anonymous):

it looks like: each term is being multiplied by 3/4.

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}8(3/4)^{n}\] so it's a geometric series

OpenStudy (anonymous):

|r| < 1 so it converges to 8/(1-3/4) = 8/(1/4) = 32